LAB: Please help with calculations:
Brand of vinegar or unknown no Trial 2 Trial Trial 2 Trial 1 1. Mass of fhask (g) 2. Mass of flask + vinegar (g) 3. Mass of vinegar (e) Q B. Analysis of Vinegar Sample 1. Buret reading of NaOH, inirial (ml.) 2. Buret reading of NaOH, final (mL) 3. Volume of NaOH used (ml.) 4. Molar concentration of NaOH (mol/L) 5. Moles of NaOH added (mol) 6. Moles of CH,COOH in vinegar (mol) 7. Mass of CH,COOH in vinegar (g) 8. Percent by mass of CH,COOH in vinegar (%) 9. Average percent by mass of CH,COOH 0 40730|1225 1340 IMI MI lllllllllll | 11 ull 11 lun 1 in vinegar (%) Calculations for Trial 1 of the first vinegar sample on next page. Experiment 10 147
Answer #1

Q5

mol fo NAOH = MV = (0.0981)(8.40*10^-3) = 0.00082404 mol of NaOH added

Q6

mol fo CH3COOH = mol of base = 0.00082404 mol

Q7

mass of CH3COO = mol*MW = 0.00082404*60 = 0.0494424 g of CH3COOH

Q8

% mass = mass of vinegar / Total mass = 0.0494424 / 1.68 *100 = 2.94%